)Consider a 아 which take on values on tix A-C,2 N)、N hare N-5c ha Joint distribution of χ anb y is Cows P(x+)
Answer #1

P(X = Y +1)

note that there are N^2 tuples of (x,y)

in this number of cases with

x = y+1

are N-1

hence

required probability

= (N - 1)/N^2